### v2

# Modified Dai-Liao Spectral Conjugate Gradient Method with Application to Signal Processing

Abstract
In this article, we present a modified variant of the Dai-Liao spectral conjugate gradient method, developed through an analysis of eigenvalues and inspired by a modified secant condition. We show that the proposed method is globally convergent for general nonlinear functions under standard assumptions. By incorporating the new secant condition and a quasi-Newton direction, we introduce updated spectral parameters. These changes ensure that the resulting search direction satisfies the sufficient descent property without relying on any line search. Numerical experiments show that the proposed algorithm performs better than several existing methods in terms of convergence speed and computational efficiency. Its effectiveness is further demonstrated through an application in signal processing.
keywords:
Spectral Conjugate Gradient Method , Dai-Liao Method , Optimization , Signal Processing
MSC:
[2020] 65K05 , 90C30 , 49M37
[…]
organization=Department of Mathematics, Banaras Hindu University,city=Varanasi, country=India
[…]
organization=Department of Mathematics, The Technion- Israel Institute of Technology,city=Haifa, country=Israel
[…]
organization=Department of Applied Mathematics and Scientific Computing, Indian Institute of Technology Roorkee,country=India
1 Introduction
[…]
find $\text{find }x^{*}\in\mathbb{R}^{m}\text{ such that }f(x^{*})=\min_{x\in\mathbb{R}^{m}}f(x).$ (1)
We denote the set of minimizers of $f$ as $\Omega_{\text{min}}(f,\mathbb{R}^{m})$.
The general iterative rule for solving (1) starts from an initial point $x_{0}\in\mathbb{R}^{m}$ and generates a sequence $\{x_{n}\}$ in $\mathbb{R}^{m}$ as follows:
$x_{n+1}=x_{n}+\alpha_{n}d_{n}\text{ for all }n\in\mathbb{N}_{0}=\mathbb{N}\cup\{0\},$ (2)
where the stepsize $\alpha_{n}$ is a positive real parameter determined via exact or inexact line search, $x_{n}$ is the previous iterative point, $x_{n+1}$ is the current point and $d_{n}$ is a suitable search direction. The class of algorithms that follow the form (2) is known as line search algorithms. For such algorithms, only the search direction $d_{n}\in\mathbb{R}^{m}$ and the stepsize $\alpha_{n}\in(0,\infty)$ are required. An appropriate descent direction $d_{n}$ must be chosen such that it satisfies the descent condition:
$\langle\nabla f(x_{n}),d_{n}\rangle<0\text{ for all }n\in\mathbb{N}_{0}.$ (3)
A common choice for a descent direction is $d_{n}=-\nabla f(x_{n})$, which reduces the general line search iteration (2) to the gradient descent (GD) method:
$x_{n+1}=x_{n}-\alpha_{n}\nabla f(x_{n})\text{ for all }n\in\mathbb{N}_{0}.$
If there exists a constant $c>0$ such that
$\langle\nabla f(x_{n}),d_{n}\rangle\leq-c\|\nabla f(x_{n})\|^{2}\text{ for all}n\in\mathbb{N}_{0},$
then the vector $d_{n}$ is said to satisfy the sufficient descent condition.
Conjugate gradient (CG) methods are among the most effective algorithms for solving large-scale unconstrained optimization problems due to their low storage requirements. These methods require only first-order derivatives, making them highly suitable for large-scale applications. The first form of the CG method was introduced by Hestenes and Stiefel [1] in 1952 for solving linear systems of equations. Later, in 1964, Fletcher and Reeves [2] modified the original idea and developed the first nonlinear CG method for unconstrained optimization problems. Due to the strong theoretical foundation and excellent numerical performance of these algorithms, many variants of the CG method have been proposed (see, e.g., [3, 4, 5, 6]).
The successive approximations generated by the CG algorithms have the form (2), where $\alpha_{n}\in(0,\infty)$ is the step length determined through a line search along the CG direction $d_{n}$. The CG directions $d_{n}$ is defined as:
$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\nabla f(x_{n})+\beta_{n-1}d_{n-1},&\text{ if }n\geq 1,\end{cases}$ (4)
where $\beta_{n}$ is the CG parameter. There are many formulas for $\beta_{n}$, each resulting in different computational behavior [7].
Based on an extended conjugacy condition, one of the fundamental CG methods was proposed by Dai and Liao [4] with the CG parameter defined as:
$\beta_{n}^{DL}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-t\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$ (5)
where $t\in(0,\infty)$, $s_{n}=x_{n+1}-x_{n}$, and $y_{n}=\nabla f(x_{n+1})-\nabla f(x_{n})$. Note that when $t=0$, $\beta_{n}^{DL}$ simplifies to the CG parameter proposed by Hestenes and Stiefel [1]:
$\beta_{n}^{HS}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
The CG parameter proposed by Hager and Zhang [8] is defined as:
$\beta_{n}^{HZ}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$
which can be viewed as an adaptive version of (5) for the specific choice $t=2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}$. Similarly, the CG parameter proposed by Dai and Kou [9] is defined as:
$\beta_{n}^{DK}(\tau_{n})=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-\left(\tau_{n}+\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},y_{n}\rangle},$
where $\tau_{n}$ is a parameter corresponding to the scaling factor in the scaled memoryless Broyden-Fletcher-Goldfarb-Shanno (BFGS) method, which can be regarded as another adaptive form of (5).
Inspired by Shanno’s matrix perspective on CG algorithms [10], Babaie-Kafaki and Ghanbari [11] observed that the Dai-Liao search directions can be expressed as $d_{n+1}=-P_{n+1}\nabla f(x_{n+1})$ for all $n\in\mathbb{N}_{0}$, where $P_{n+1}=I-\frac{s_{n}y_{n}^{T}}{\langle s_{n},y_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}$ is a nonsingular matrix for $t\in(0,\infty)$ and $\langle s_{n},y_{n}\rangle\neq 0$; see [12]. Conducting an eigenvalue analysis on a symmetrized version of $P_{n+1}$, defined by $A_{n+1}=\frac{P_{n+1}+P_{n+1}^{T}}{2}$, we can derive a two-parameter formula for $t$ in the Dai-Liao CG parameter (5), given by
$t_{n}^{p,q}=p\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}-q\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0},$ (6)
which ensures the descent condition (3) for $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. However, in cases where $\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}$ and $\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$ are nearly equal, their subtraction may not be numerically stable [13]. Moreover, the parameter $t$ in CG parameter $\beta_{n}^{DL}$ must be nonnegative. To address these issues, they restricted $q$ in (6) to be nonpositive and proposed the following CG parameter:
$\beta_{n}^{DDL}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$
where $t_{n}^{p,q}$ is defined by (6) with $p\in\left(\frac{1}{4},\infty\right)$ and $q\in(-\infty,0]$. Note that for the parameter choice $(p,q)=(2,0)$ in (6), $\beta_{n}^{DDL}$ reduces to $\beta_{n}^{HZ}$. Similarly, if $(p,q)=(1,0)$, then $\beta_{n}^{DDL}$ is equivalent to $\beta_{n}^{DK}(\tau_{n})$ with the optimal choice of $\tau_{n}=\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$.
Another significant category of iterative methods used for solving unconstrained optimization problems is the spectral gradient method. Initially introduced by Barzilai and Borwein [14], these methods were later advanced by Raydan [15], who developed a spectral gradient approach specifically for large-scale unconstrained optimization challenges. The spectral conjugate direction is defined by
$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\theta_{n}\nabla f(x_{n})+\beta_{n-1}d_{n-1},&\text{ if }n\geq 1,\end{cases}$
where $\{\theta_{n}\}\subset(0,\infty)$ is spectral parameter and $\{\beta_{n}\}\subset(0,\infty)$ is conjugate parameter.
Building on the encouraging numerical results associated with spectral gradient methods [14, 15], Birgin and Martínez [16] proposed the integration of spectral gradient techniques with conjugate gradient methods, culminating in the development of the first spectral conjugate gradient (CG) method. The spectral CG methods represent a powerful enhancement of traditional CG methods by incorporating the principles of spectral gradients within the conjugate gradient framework. Furthering this line of research, Yu et al. [17] modified the spectral Perry’s CG method [18], resulting in an efficient spectral CG variant. Capitalizing on the robust convergence properties of quasi-Newton methods, Andrei [19] introduced a scaled BFGS preconditioned CG method in 2007, leveraging quasi-Newton techniques to enhance the effectiveness of conjugate gradient methods. This foundational idea has been elaborated upon by Andrei [20, 21, 22], as well as Babaie-Kafaki and Mahdavi-Amiri [23].
Using the modified secant condition proposed by Li and Fukushima [24], Zhou and Zhang [25] established the following modified secant condition:
$\nabla^{2}f(x_{n+1})s_{n}=z_{n},$
where
$z_{n}=y_{n}+h_{n}\|\nabla f(x_{n})\|^{r}s_{n},\quad h_{n}=\nu+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|\nabla f(x_{n})\|^{-r}$ (7)
with $r,\nu\in(0,\infty)$. The modified secant condition (7) plays a crucial role in establishing the global convergence of the MBFGS method for non-convex functions; see [25].
It is worth noting that the following relation holds independently of the line search:
$\langle z_{n},s_{n}\rangle$ $=\langle y_{n},s_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|^{2}+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|s_{n}\|^{2}$
$\geq\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|^{2}>0\text{ for all }n\in\mathbb{N}_{0}.$
Inspired from the appropriate theoretical properties given by (7), Faramarzi and Amini [26] modified the spectral CG method proposed by Jian et al. [27] and considered the following conjugate parameter
$\beta_{n}^{ZDK}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
Based on this conjugate parameter, they construct the following spectral parameter:
$\theta_{n+1}=\begin{cases}1-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle\nabla f(x_{n+1}),z_{n}\rangle},&\text{ if }1-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle\nabla f(x_{n+1}),z_{n}\rangle}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise },\end{cases}$
where $\eta$ and $\tau$ are positive constants. To ensure the global convergence of this method, Faramarzi and Amini [26] used the line search satisfying the strong Wolfe conditions:
$f(x_{n}+\alpha_{n}d_{n})\leq f(x_{n})+\delta\alpha_{n}\langle\nabla f(x_{n}),d_{n}\rangle,$ (8)
$|\langle\nabla f(x_{n}+\alpha_{n}d_{n}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle,$ (9)
where $10<\delta<\sigma<$ 1 and the following assumptions:
  1. (A1)
    The level set
    $C_{0}=\{x\in\mathbb{R}^{n}:f(x)\leq f(x_{0})\}$
    is bounded; namely, there exists a constant $M>0$ such that
    $\|x\|\leq M\text{ for all }x\in C_{0}.$
  2. (A2)
    In some neighbourhood $\Omega$ of $C_{0}$, $f$ is continuously differentiable and its gradient is Lipschitz continuous, namely there exists a constant $L>0$ such that
    $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|\text{ for all }x,y\in\Omega.$ (10)
The primary motivation of this study is to introduce a modified spectral conjugate gradient method and to establish its global convergence for general functions under standard assumptions. This new method consistently satisfies the sufficient descent property, regardless of the line search used. To demonstrate the effectiveness of the proposed approach, we present numerical results that validate its performance and show its superior convergence compared to the modified spectral conjugate gradient method of Faramarzi and Amini [26] and the scaled conjugate gradient method of Mard and Fakhari [28]. Furthermore, we explore its application in compressed sensing, emphasizing its practical relevance and significant advantages.
2 Preliminaries
In this section, we present some fundamental results from Euclidean space that will be used to establish the convergence of the proposed method.
Lemma 1.
([26]) Let $a,b,\sigma\in\mathbb{R}$. Then
$(c+\sigma d)^{2}\leq(1+\sigma^{2})(c^{2}+d^{2})\text{ for all }c,d,\sigma\geq 0,$
Lemma 2.
([26]) Let $\mathbb{R}^{m}$ be the Euclidean space. Then
$\langle u,v\rangle\leq\frac{1}{2}(\|u\|^{2}+\|v\|^{2})\text{ for all }u,v\in\mathbb{R}^{m}.$
Lemma 3.
Let $\mathbb{R}^{m}$ be the Euclidean space and let $\{u_{n}\}$ and $\{v_{n}\}$ be sequences in $\mathbb{R}^{m}$. Let $b_{0},\sigma\in(0,\infty)$ and $u_{0}\in\mathbb{R}^{m}$ and let $\{a_{n}\}$ and $\{b_{n}\}$ be sequences in $(0,\infty)$ satisfying the following conditions:
  1. (C1)
    $a_{n}>a$ for all $n\in\mathbb{N}$,
  2. (C2)
    $\langle v_{n+1},u_{n+1}\rangle<0$ for all $n\in\mathbb{N}$,
  3. (C3)
    $|\langle v_{n+1},u_{n}\rangle|\leq-\sigma\langle v_{n},u_{n}\rangle$ for all $n\in\mathbb{N}$.
Suppose that
$u_{n+1}=-a_{n+1}v_{n+1}+b_{n}u_{n}\text{ for all }n\in\mathbb{N}.$ (11)
Then
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}+a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\geq\frac{a}{1+\sigma^{2}}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$
Proof.
From (11) and condition (C2), we obtain
$b_{n}^{2}\|u_{n}\|^{2}$ $=\|u_{n+1}\|^{2}+2a_{n+1}\langle v_{n+1},u_{n+1}\rangle+a_{n+1}^{2}\|v_{n+1}\|^{2}$
$\leq\|u_{n+1}\|^{2}+a_{n+1}^{2}\|v_{n+1}\|^{2},$
which implies that
$\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\geq b_{n}^{2}-a^{2}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$ (12)
Again, from (11), we have
$-\langle v_{n+1},u_{n+1}\rangle+b_{n}\langle v_{n+1},u_{n}\rangle=a_{n+1}\|v_{n+1}\|^{2}.$ (13)
From condition (C3), we have
$b_{n}\langle v_{n+1},u_{n}\rangle\leq b_{n}|\langle v_{n+1},u_{n}\rangle|\leq-\sigma b_{n}\langle v_{n},u_{n}\rangle\leq\sigma b_{n}|\langle v_{n},u_{n}\rangle|.$ (14)
From (13) and (14), we get
$-\langle v_{n+1},u_{n+1}\rangle+\sigma b_{n}|\langle u_{n},v_{n}\rangle|\geq a_{n+1}\|v_{n+1}\|^{2}.$
Squaring both sides and applying Lemma 1 with $c=-\langle v_{n+1},u_{n+1}\rangle>0$ (from (C2)) and $d=b_{n}|\langle v_{n},u_{n}\rangle|>0$ gives
$\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}\geq\frac{a_{n+1}^{2}}{1+\sigma^{2}}\|v_{n+1}\|^{4}.$
Since $a_{n+1}>a$, this yields that
$\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}$ $\geq\frac{a}{1+\sigma^{2}}\|v_{n+1}\|^{4},$
which implies that
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}$
$=\frac{1}{\|u_{n+1}\|^{2}}\left[\langle v_{n+1},u_{n+1}\rangle^{2}+\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\langle v_{n},u_{n}\rangle^{2}\right]$
$\geq\frac{1}{\|u_{n+1}\|^{2}}\left[\frac{a}{1+\sigma^{2}}\|v_{n+1}\|^{4}+\left(\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}-b_{n}^{2}\right)\langle v_{n},u_{n}\rangle^{2}\right].$ (15)
From (12) and (15), we conclude that
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n+1},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}\geq\frac{\|v_{n+1}\|^{4}}{\|u_{n+1}\|^{2}}\left[\frac{a}{1+\sigma^{2}}-a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\right],$
which implies that
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}+a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\geq\frac{a}{1+\sigma^{2}}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$
Lemma 4.
([29]) Let $f:\mathbb{R}^{n}\to\mathbb{R}$ such that assumptions (A1) and (A2) hold. Consider any iterative method of the form (2), where $d_{n}$ is a descent direction and $\alpha_{n}$ is obtained by the strong Wolfe conditions (8) and (9). Then
$\sum_{n=0}^{\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}<\infty.$
From Lemma 4, we have
$\lim_{n\to\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}=0.$ (16)
3 A modified descent Dai-Liao spectral conjugate gradient method and its convergence analysis
In this section, we introduce a modified spectral conjugate gradient (CG) method, inspired by ideas from quasi-Newton methods.
Using the theoretical properties provided by (7) and the Dai-Liao conjugate gradient parameter (5), we propose the following modified Dai-Liao conjugate parameter:
$\beta_{n}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$ (17)
where $t$ is a nonnegative parameter.
Note that, by (4) and (17), the search direction can be expressed as
$d_{n+1}=-Q_{n+1}\nabla f(x_{n+1})\text{ for all }n\in\mathbb{N}_{0},$
where
$Q_{n+1}=I-\frac{s_{n}z_{n}^{T}}{\langle s_{n},z_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}.$
Following the methodology in [11], we propose the following formula for the parameter $t$ in the modified Dai-Liao conjugate parameter (17):
$t_{n}^{p,q}=p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0}$ (18)
with $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. Using this parameter $t_{n}^{p,q}$, we define the following modified descent Dai-Liao (MDDL) conjugate parameter:
$\beta_{n}^{MDDL}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$ (19)
Now, we consider the following spectral conjugate gradient search direction:
$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\theta_{n}\nabla f(x_{n})+\beta_{n-1}^{MDDL}d_{n-1}&\text{ if }n\geq 1,\end{cases}$ (20)
where $\{\theta_{n}\}$ is a sequence in $(0,\infty)$.
Proposition 5.
Let $f:\mathbb{R}^{m}\to\mathbb{R}$ be a differentiable function and let $\{d_{n}\}$ be a sequence of spectral CG search directions generated by (20). Then, we have the following:
  • (a)
    For all $n\in\mathbb{N}_{0},$ we have
    $\langle\nabla f(x_{n}),d_{n}\rangle\leq-\kappa_{n}\|\nabla f(x_{n})\|^{2},$ (21)
    where
    $\kappa_{n}=\left\{\begin{array}[]{ll}1,&\text{if }n=0,\\ \theta_{n}-\left(\frac{1}{4p}+|q|\right),&\text{if }n\geq 1.\end{array}\right.$
  • (b)
    If $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$, then the direction given by (20) is a descent direction.
Proof.
(a) Note that $\langle\nabla f(x_{0}),d_{0}\rangle=-\|\nabla f(x_{0})\|^{2}$. Then the inequality (21) holds for $n=0$. We show that the inequality (21) holds for $n>0$. Let $n\in\mathbb{N}_{0}$. Note that $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. From (18), (19) and (20), we get
$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $=\langle\nabla f(x_{n+1}),-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\rangle$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-\frac{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-\left(p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}.$
Since $s_{n}=\alpha_{n}d_{n}$, we have
$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}+q\frac{\langle d_{n},z_{n}\rangle}{\|d_{n}\|^{2}}\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}+q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}.$ (22)
From Lemma 2 with $u_{n}=\frac{1}{\sqrt{2p}}\nabla f(x_{n+1})$ and $v_{n}=\frac{\sqrt{2p}\langle\nabla f(x_{n+1}),d_{n}\rangle z_{n}}{\langle d_{n},z_{n}\rangle},$ we obtain
$\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\leq\frac{1}{4p}\|\nabla f(x_{n+1})\|^{2}+\frac{p\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}.$ (23)
On the other hand, by Cauchy-Schwarz inequality, we have
$q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}\leq|q|\frac{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}}{\|d_{n}\|^{2}}=|q|\|\nabla f(x_{n+1})\|^{2}.$ (24)
Combining (22), (23) and (24), we conclude that
$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $\leq-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{1}{4p}\|\nabla f(x_{n+1})\|^{2}+|q|\|\nabla f(x_{n+1})\|^{2}$
$=-\left(\theta_{n+1}-\left(\frac{1}{4p}+|q|\right)\right)\|\nabla f(x_{n+1})\|^{2}.$
(b) Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$. Then, from (21), we get $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Therefore, $\{d_{n}\}$ is a descent direction. ∎
Now, we construct the spectral parameter for our conjugate gradient method based on the conjugate parameter $\beta_{n}^{MDDL}$.
Quasi-Newton methods are a prominent class of iterative optimization techniques that do not require computing the exact Hessian matrix at each iteration. The primary motivation behind these methods is to achieve the rapid convergence rate of the classical Newton method while reducing the computational cost. Instead of calculating the Hessian matrix directly, quasi-Newton methods construct a sequence of Hessian approximations. Inspired by the quasi-Newton direction, the search direction $d_{n+1}$ defined by (20) is designed to mimic the behavior of the quasi-Newton method. Consequently, we aim to determine a suitable spectral parameter $\theta_{n+1}$ such that
$-B_{n+1}^{-1}\nabla f(x_{n+1})\approx-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n},$
where $B_{n+1}$ is an approximation of the Hessian matrix $\nabla^{2}f(x_{n+1})$. Taking inner product with $B_{n+1}s_{n}$ gives that
$-\langle s_{n},\nabla f(x_{n+1})\rangle\approx-\theta_{n+1}\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle,$
which implies that
$\theta_{n+1}\approx\frac{1}{\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle).$
Since $B_{n+1}$ must satisfies the modified secant condition $B_{n+1}s_{n}=z_{n}$, we have
${\theta_{n+1}}$ $=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle z_{n},d_{n}\rangle)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\left(\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right.\right.$
$\quad\left.\left.-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)\langle z_{n},d_{n}\rangle\right)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right.$
$\quad\left.-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)$
$=\frac{\langle s_{n},\nabla f(x_{n+1})\rangle+\langle\nabla f(x_{n+1}),z_{n}\rangle-t_{n}^{p,q}\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}$
$=1-\frac{{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}-\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}$ (25)
$=1-(t_{n}^{p,q}-1)\frac{\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}.$ (26)
We neglect $\langle s_{n},\nabla f(x_{n+1})\rangle$ in (25), which leads to
${\theta_{n+1}}=1-\frac{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}.$ (27)
We denote the $\theta_{n+1}$ obtained in equation (26) by $\theta_{n+1}^{R}$ and the $\theta_{n+1}$ obtained in equation (27) by $\theta_{n+1}^{N}$. To obtain the sufficient descent condition and the bounded property of the spectral parameters, we propose the following spectral parameters:
$\theta_{n+1}=\begin{cases}\theta_{n+1}^{R},&\text{ if }\theta_{n+1}^{R}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise }\end{cases}$ (28)
or
$\theta_{n+1}=\begin{cases}\theta_{n+1}^{N},&\text{ if }\theta_{n+1}^{N}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise },\end{cases}$ (29)
where $\eta$ and $\tau$ are positive constants. Relation (28) and (29) lead to
$\frac{1}{4p}+|q|+\eta\leq\theta_{n+1}\leq\tau\textrm{ for all }n\in\mathbb{N}_{0}.$ (30)
Now, we propose the following modified descent Dai-Liao spectral conjugate gradient method (MDDLSCG):
Input: Let $x_{0}\in\mathbb{R}^{n}$ and let $\epsilon,\delta,\sigma,\eta,\tau,r,\nu\in(0,\infty)$ such that
     $10<\delta<\sigma<$ 1.
Step 0: Set $d_{0}=-\nabla f(x_{0})$.
Step 1: If $\|\nabla f(x_{n})\|_{\infty}<\epsilon$, stop.
Step 2: Determine the step length $\alpha_{n}$ satisfying the strong Wolfe
    conditions (8) and (9).
Step 3: Set $x_{n+1}=x_{n}+\alpha_{n}d_{n}$.
Step 4: Compute $d_{n+1}$ by
$d_{n+1}=-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}d_{n}\text{ for all }n\in\mathbb{N}_{0},$
      where CG parameter $\beta_{n}^{MDDL}$ is given by (19) and spectral
    parameter is given by (28) or (29).
Step 5: Set $n=n+1$ and return to Step 1.
Algorithm 1 MDDLSCG Algorithm
Remark 6.
(i) Using $\theta_{n}\geq\frac{1}{4p}+|q|+\eta$ and (21), we conclude that
$\langle\nabla f(x_{n}),d_{n}\rangle\leq-\eta\|\nabla f(x_{n})\|^{2}\text{ for all }n\in\mathbb{N}.$ (31)
(ii) If $\|\nabla f(x_{n})\|\neq 0$, then, from (9) and (31), we see that
$\langle d_{n},y_{n}\rangle\geq-(1-\sigma)\langle\nabla f(x_{n}),d_{n}\rangle\geq(1-\sigma)\eta\|\nabla f(x_{n})\|^{2}>0,$ (32)
which, from (7), implies that $h_{n}=\nu$ and hence
$z_{n}=y_{n}+\nu\|\nabla f(x_{n})\|^{r}s_{n}.$ (33)
Now, we analyze the global convergence of the MDDLSCG Algorithm.
Remark 7.
  1. (i)
    From (8) and (31), we can see that $\{f(x_{n})\}$ is decreasing. Thus, the sequence $\{x_{n}\}$ generated by Algorithm 1 is contained in $C_{0}$.
  2. (ii)
    From Assumption (A2), we have
    $\|\nabla f(x)\|$ $\leq\|\nabla f(x)-\nabla f(x_{0})\|+\|\nabla f(x_{0})\|$
    $\leq L\|x-x_{0}\|+\|\nabla f(x_{0})\|$
    $\leq 2ML+\|\nabla f(x_{0})\|:=\gamma\text{ for all }x\in C_{0}.$ (34)
In addition, suppose that $\nabla f(x_{n})\neq 0$ for all $n\in\mathbb{N}_{0}$; otherwise, a stationary point will be found.
Lemma 8.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$. If $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$, then $\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}}<\infty.$
Proof.
Since $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$. Then, there exists $\epsilon>0$ such that $\|\nabla f(x_{n})\|\geq\epsilon$ for all $n\in\mathbb{N}_{0}$. This together with (16) implies that
$\lim_{n\to\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}}=0.$ (35)
Let $u_{n}=d_{n}$, $v_{n}=\nabla f(x_{n})$, $a_{n}=\theta_{n}$ and $b_{n}=\beta_{n}$ in Lemma 3. Since $\theta_{n}>\frac{1}{4p}+|q|$, it follows from Remark 6 that $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Also from strong Wolfe condition (9), we have $|\langle\nabla f(x_{n+1}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle$. Then, clearly all the conditions (C1), (C2) and (C3) of Lemma 3 are satisfied. Thus, we have
$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq b_{1}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}-\theta_{n+1}^{2}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}},$
where $b_{1}=\left(\frac{1}{4p}+|q|\right)\frac{1}{1+\sigma^{2}}$. It follows from (35) that for all sufficiently large $n$, we have
$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq b_{1}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}.$ (36)
From (36) and Lemma 4, we get that $\sum_{n=1}^{\infty}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}<\infty.$
Theorem 9.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Then $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0.$
Proof.
Suppose the claim of the theorem is not true, i.e., there exists a positive constant $\xi$ such that
$\|\nabla f(x_{n})\|>\xi\text{ for all }n\in\mathbb{N}_{0}.$ (37)
From (32), (33) and (37), we obtain
$\langle d_{n},z_{n}\rangle$ $=\langle d_{n},y_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}$
$\geq\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}\geq\nu\xi^{r}\alpha_{n}\|d_{n}\|^{2}.$ (38)
From (33), (10) and ((ii)), we have
$\|z_{n}\|\leq\|y_{n}\|+\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|\leq L\|s_{n}\|+\nu\gamma^{r}\|s_{n}\|=(L+\nu\gamma^{r})\|s_{n}\|.$ (39)
From the Lipschitz continuity of the $\nabla f$ and (39), we obtain
$t_{n}^{p,q}$ $\leq\left|p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right|$
$\leq p\frac{(L+\nu\gamma^{r})^{2}\|s_{n}\|^{2}}{\mu\|s_{n}\|^{2}}+|q|\frac{(L+\nu\gamma^{r})\|s_{n}\|^{2}}{\|s_{n}\|^{2}}=p\frac{L^{2}}{\mu}+L|q|.$
Thus, the relations (19), (10), ((ii)), (38), (39), along with Cauchy-Schwarz inequality, conclude
$\|\beta_{n}^{MDDL}\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\left(p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|+p\left\|\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$\quad+|q|\left\|\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\frac{\|\nabla f(x_{n+1})\|\|z_{n}\|}{\langle d_{n},z_{n}\rangle}+p\frac{\|z_{n}\|^{2}\|\nabla f(x_{n+1})\|\|d_{n}\|}{\langle d_{n},z_{n}\rangle^{2}}+|q|\frac{\|\nabla f(x_{n+1})\|}{\|d_{n}\|}$ (40)
$\leq\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}\|d_{n}\|}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}\|d_{n}\|}+|q|\frac{\gamma}{\|d_{n}\|}.$
Note that $\theta_{n+1}\leq\tau$ by (30). Hence, from (20), (40), we get
$\|d_{n+1}\|$ $=\|-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\|$
$\leq\|\theta_{n+1}\|\|\nabla f(x_{n+1})\|+\|\beta_{n}^{MDDL}\|\|d_{n}\|$
$\leq\tau\gamma+\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}}+|q|\gamma.$ (41)
From (41), we observe that
$\frac{1}{\|d_{n+1}\|^{2}}\geq\frac{1}{E^{2}}\textrm{ for all }n\in\mathbb{N}_{0},$
where $E=\tau\gamma+\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}}+|q|\gamma$. Thus,
$\sum_{n=0}^{\infty}\frac{1}{\|d_{n}\|^{2}}=\infty.$ (42)
From (37) and Lemma 8, we have
$\sum_{n=0}^{\infty}\frac{1}{\|d_{n}\|^{2}}=\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}\|\nabla f(x_{n})\|^{4}}\leq\frac{1}{\xi^{4}}\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}}<\infty,$
which contradicts (42). Therefore, $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0$. ∎
4 Numerical Experiments
In this section, we discuss the numerical performance of our proposed method (MDDLSCG), comparing it with Algorithm 3.1 in [26] (denoted by MSCG) and Scaled conjugate gradient method (denoted by ScCG) in [28]. All numerical experiments were conducted in MATLAB R2024b on an HP workstation with Intel Core i7-13700 processor, 32 GB RAM and Windows 11.
For the numerical experiment, we consider two test problems: (i) finding the minima of the Beale function, and (ii) finding the minima of the Almost Perturbed Quadratic (APQ) function:
Beale Test Function: The Beale function is multimodal with sharp peaks at the corners of the input domain, which is defined from the general Beale function as follows:
$f(x,y)=\left(1.5-x+xy\right)^{2}+\left(2.25-x+xy^{2}\right)^{2}+\left(2.625-x+xy^{3}\right)^{2}\text{ for all }(x,y)\in\mathbb{R}^{2}.$
The global minimum is located at $(x,y)=(3,0.5)$, where $f(x,y)=0$.
During the numerical experiments, we set the parameters $\delta=0.01$, $\sigma\in\{0.1,0.2,0.4,0.6,0.8,0.9\}$ for all the algorithms, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$ for algorithms MSCG and MDDLSCG and $p=0.4$ and $q=0.2$ for the algorithm MDDLSCG. We take the initial point $x_{0}=(1,0.8)$. During the implementation, the iterations are terminated, when it meets the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-15}$.
The numerical results of all the algorithms MDDLSCG, MSCG and ScCG for the Beale function are summarized in Table 1, while their performance is illustrated in Figure LABEL:bealefig1 and LABEL:bealefig2. In Figures LABEL:bealefig1 (a), (c), (e), the iteration paths are visualized on the contour plot of the Beale function, showing how each method converges to the minimizer for $\sigma=0.1,0.2,0.6$, respectively. Figures LABEL:bealefig1 (b), (d), (f) depict the convergence rate by plotting the norm of the gradient, $\|\nabla f(x_{n})\|_{\infty}$ versus the number of iterations on a logarithmic scale for $\sigma=0.1,0.2,0.6$, respectively. The graph between different values of sigma and number of iterations is given in Figure LABEL:bealefig2 (a) and the graph between different values of sigma and execution time is in Figure LABEL:bealefig2 (b).
From the data presented, it is evident that our proposed method, MDDLSCG, outperforms both MSCG and ScCG in all tested scenarios. As shown in Table 1, Figure LABEL:bealefig1 and LABEL:bealefig2 (b), MDDLSCG reaches the stopping criteria in fewer iterations and requires less execution time. Additionally, From Figure LABEL:bealefig2 (a), we observe that these methods are significantly sensitive to variations in the parameter $\sigma$, which affects their performance.
$\sigma$ MDDLSCG MSCG ScCG itr Tcpu $E_{n}$ itr Tcpu $E_{n}$ itr Tcpu $E_{n}$ 0.1 37 1.86e-02 9.49e-15 63 3.14e-02 2.63e-16 200 5.40e-02 5.27e-16 0.2 36 1.75e-02 3.16e-15 63 2.76e-02 9.07e-15 136 3.99e-02 4.51e-15 0.4 53 2.44e-02 9.40e-15 73 2.65e-02 8.75e-15 126 4.73e-02 2.87e-15 0.6 40 2.36e-02 7.91e-16 72 2.78e-02 5.74e-15 95 4.94e-02 9.32e-15 0.8 67 3.24e-02 5.32e-15 111 2.86e-02 4.48e-15 201 4.92e-02 6.91e-15 0.9 64 3.28e-02 8.16e-15 101 3.48e-02 8.74e-15 184 4.37e-02 1.08e-15
Table 1: Comparison table for MDDLSCG, MSCG and ScCG for Beale function
Almost Perturbed Quadratic function: Consider the Almost Perturbed Quadratic (APQ) function $f:\mathbb{R}^{m}\to\mathbb{R}$ defined by
$f(\mathbf{x})=\frac{1}{2}\mathbf{x}^{T}A\mathbf{x}+\mathbf{b}^{T}\mathbf{x}$
where $A\in\mathbb{R}^{m\times m}$ is a symmetric positive definite matrix ans $b\in\mathbb{R}^{m}$ is a vector of perturbation. This function has a global minimum at $x=-A^{-1}b$. Note that the gradient of $f$ is given by
$\nabla f(x)=Ax+b\text{ for all }x\in\mathbb{R}^{m}.$
For the numerical experiments, we have considered three cases, each characterized by dimensions of1000100010001000, $100$ and $25$. To the cases with dimensions $1000$ and $100$, the matrix $A$ and vector $b$ are randomly generated using MATLAB’s built-in function rand and the initial point is randomly generated using MATLAB’s rand function.
For the case with dimension $25$, we take the initial point
$x_{0}$ $=(x_{0,1},\ldots,x_{0,i},\ldots,x_{0,25}),\text{ where }x_{0,i}=1\text{ for all }i=1,2,\ldots,25\text{ and}$
$a$ $=(2,4,3,6,4,12,8,10,5,8,6,11,4,8,3,6,2,12,6,8,4,10,5,2,6),$
$b$ =(4,52,18,60,24,72,72,50,55,80,78,22,12=(-4,-52,-18,-60,-24,-72,-72,-50,-55,-80,-78,-22,-12,= ( - 4 , - 52 , - 18 , - 60 , - 24 , - 72 , - 72 , - 50 , - 55 , - 80 , - 78 , - 22 , - 12 ,
80,27,42,24,120,30,120,24,100,60,20,66)$\quad$-80,-27,-42,-24,-120,-30,-120,-24,-100,-60,-20,-66),- 80 , - 27 , - 42 , - 24 , - 120 , - 30 , - 120 , - 24 , - 100 , - 60 , - 20 , - 66 ) ,
$A$ $=\text{diag}(a).$
For all the cases, the parameters are set as follows: $\delta=0.01$ and $\sigma=0.1$ for all algorithms, while for MSCG and MDDLSCG, we use $\eta=0.001$, $\tau=10$, $r=1$, and $\nu=0.001$. Additionally, for the MDDLSCG algorithm, we set $p=0.4$ and $q=0.2$.
During the implementation, the iterations are terminated when the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-6}$ is satisfied, where $x_{n}=(x_{n,1},\ldots,x_{n,25})$.
The performance of the algorithms MDDLSCG, MSCG, and ScCG for the APQ function is summarized in Table 2 for all three cases. To provide a clearer visualization of the experiment, we have plotted the graph of the number of iterations versus $E_{n}$ in Figure LABEL:APQgraph. From Table 2 and Figure LABEL:APQgraph, it is evident that our proposed method, MDDLSCG, requires fewer iterations and less computational time to satisfy the desired stopping criterion across all cases.
Figure LABEL:APQtrack illustrates how the algorithms MDDLSCG, MSCG, and ScCG approximate the desired solution $x^{*}=-A^{-1}b$ at the 5th and 10th iterations for the case where the problem dimension is $25$.
Furthermore, Figure LABEL:APQprogress depicts the convergence behavior of the algorithms as the number of iterations increases up to the 50th iteration for the same dimension. It is observed that MDDLSCG demonstrates superior convergence characteristics compared to MSCG and ScCG.
Dimension Method itr Tcpu $Er_{n}$
25 MDDLSCG 52 3.088055e-02 5.691882e-07
MSCG 59 3.947327e-02 6.931169e-07
ScCG 65 4.730704e-02 4.514777e-07
100 MDDLSCG 102 5.552326e-02 8.162825e-07
MSCG 121 6.512598e-02 9.411660e-07
ScCG 134 6.823947e-02 8.609522e-07
1000 MDDLSCG 132 3.481332e-02 9.291368e-07
MSCG 146 3.522395e-02 9.484382e-07
ScCG 171 4.053286e-02 8.747078e-07
Table 2: Comparison table of MDDLSCG, MSCG and ScCG for APQ function
5 Application in compressed sensing
Compressed sensing is a well-known signal processing strategy for efficiently acquiring and reconstructing signals with a sparse structure, allowing for memoryless and compact storage of the signals [30]. The importance of compressed sensing in real-world applications, such as machine learning, compressive imaging, radar, wireless sensor networks, medical imaging, astrophysical signals, and video coding, has been highlighted in [31].
The compressed sensing problem involves finding a sparse solution to a highly underdetermined system, represented as $Ax=b$, where $A\in\mathbb{R}^{m\times n}$ (with $m\ll n$) and $b\in\mathbb{R}^{m}$ [30]. A common approach is to solve the following optimization problem:
$f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\phi(x),$
where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a penalty function, and $\mu>0$ is a parameter that controls the trade-off between sparsity and reconstruction fidelity. A popular choice for $\phi$ is $\phi(x)=\|x\|_{1}$, leading to the well-known Basis Pursuit Denoising (BPD) problem, which has been extensively studied in the literature; see [32] and the references therein. However, solving BPD is challenging due to the non-smoothness of the $\ell_{1}$-penalty term. To address this issue, Zhu et al. [33] proposed a relaxation of BPD using Nesterov’s smoothing technique [34], resulting in the following formulation:
$f_{\lambda}(x)=\sum_{i=1}^{n}\psi_{\lambda}(|x_{i}|),$
where $x_{i}$ represents the $i$-th component of the vector $x$, and $\psi_{\lambda}$ is defined as:
$\psi_{\lambda}(|x_{i}|)=\max_{\gamma\in[0,1]}\left\{\gamma|x_{i}|-\frac{1}{2}\lambda\gamma^{2}\right\},$
which simplifies to:
$\psi_{\lambda}(|x_{i}|)=\begin{cases}\frac{|x_{i}|^{2}}{2\lambda},&\text{if }|x_{i}|<\lambda,\\ |x_{i}|-\frac{\lambda}{2},&\text{if }|x_{i}|\geq\lambda.\end{cases}$
Here, $f_{\lambda}$ is known as the Huber function [35], and its gradient $\nabla f_{\lambda}$ has a Lipschitz constant of $\frac{1}{\lambda}$ [36], with the explicit form given by:
$(\nabla f_{\lambda})_{i}=\begin{cases}\frac{x_{i}}{\lambda},&\text{if }|x_{i}|<\lambda,\\ \text{sgn}(x_{i}),&\text{otherwise},\end{cases}$
where $\lambda>0$ is a constant and $\text{sgn}(\cdot)$ denotes the signum function.
Therefore, instead of solving the original problem $f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\|x\|_{1}$, we solve the smoothed version:
$\min_{x\in\mathbb{R}^{n}}f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu f_{\lambda}(x).$
Clearly, the problem above is an unconstrained smooth convex optimization problem, and its gradient is given by:
$\nabla f(x)=\lambda\nabla f_{\lambda}(x)+A^{T}(Ax-b).$
For the numerical experiments, we aim to reconstruct a sparse signal of length $n$ from $m$-length observations ($m\ll n$). The parameters for the MDDLSCG, MSCG and ScCG algorithms are set as follows:
$\mu=\max\{2^{-7},0.001\|A^{T}b\|_{\infty}\}$, $\lambda=\min\{0.001,0.048\|A^{T}b\|_{\infty}\}$, $\sigma=0.1$, $\delta=0.01$, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$.
[…]
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-x_{i}^{*})^{2}\leq 10^{-5}.$
To evaluate the restoration performance qualitatively, we report the relative error (RelErr) of the recovered signal, which is defined as follows:
$\text{RelErr}=\frac{\|x^{*}-x\|}{\|x\|}.$
Let $k$ denote the number of nonzero elements in the original signal. For a fair comparison, we consider two different settings: $(m,n,k)=(128,512,16)$, $(256,1024,32)$.
The detailed numerical results are presented in Table 3, where we report the average number of iterations (Itr), average CPU time (Tcpu), mean squared error (MSE), and relative error (RelErr) for each setting. The graph depicting the relationship between relative error and the number of iterations is shown in Figures LABEL:stpp512(a) and LABEL:stopp(a), while the graph illustrating MSE versus the number of iterations is presented in Figures LABEL:stpp512(b) and LABEL:stopp(b). Additionally, the recovery of the signal is visualized in Figures 11 and 13.
From Table 3, Figures LABEL:stpp512, and LABEL:stopp, it is clear that the MDDLSCG algorithm surpasses the MSCG and ScCG algorithms in performance across all cases.
Figure 11: Signal recovery for data $(m,n,k)=(128,512,16)$
Figure 13: Signal recovery for data $(m,n,k)=(256,1024,32)$
Dim$(m,n,k)$ Method Itr Tcpu MSE RelErr (128,512,16) MDDLSCG 272 1.186382e-01 2.033486e-06 7.487872e-03 MSCG 493 1.606987e-01 2.033495e-06 7.488087e-03 ScCG 3653 1.135658e+00 2.033588e-06 7.712389e-03 (256,1024,32) MDDLSCG 291 2.163503e-01 2.128598e-06 9.721181e-03 MSCG 1405 8.339790e-01 2.128516e-06 9.720993e-03 ScCG 3002 1.746555e+00 2.229608e-04 9.949161e-02
Table 3: Comparison table for methods MDDLSCG, MSCG and ScCG
6 Conclusions
We investigated a modified descent Dai-Liao spectral conjugate gradient (MDDLSCG) method that incorporates a new secant condition and quasi-Newton direction. The updated method introduces spectral parameters ensuring a sufficient descent property independent of line search techniques. Theoretical analysis establishes the global convergence of the method for general nonlinear functions under standard assumptions. Numerical experiments and an application in signal processing demonstrate the improved performance and practical effectiveness of the proposed algorithm. Additionally, we present numerical results that validate its performance and show its superior convergence compared to the MSCG method [26] and the ScCG method [28], highlighting its potential in optimization and related applications.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
The authors are grateful to the editor and to the reviewers for their remarks and advice which improved our manuscript.
Funding
The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.