### v3

# A Modified Dai-Liao Spectral Conjugate Gradient Method with an Application to Signal Processing

Abstract
We propose and study a variant of the Dai-Liao spectral conjugate gradient method, developed through an analysis of eigenvalues and inspired by a modified secant condition. We show that our proposed method is globally convergent for general nonlinear functions under standard assumptions. By incorporating the new secant condition and a quasi-Newton direction, we introduce updated spectral parameters. These changes ensure that the resulting search direction satisfies the sufficient descent property without relying on any line search. Numerical experiments show that the proposed algorithm performs better than several existing methods in terms of convergence speed and computational efficiency. Its effectiveness is further demonstrated through an application to signal processing.
keywords:
Dai-Liao Method , Optimization , Signal Processing , Spectral Conjugate Gradient Method
2020 MSC:
65K05 , 90C30 , 49M37
[…]
organization=Department of Mathematics, Banaras Hindu University,city=Varanasi, country=India
[…]
organization=Department of Mathematics, The Technion – Israel Institute of Technology,city=Haifa, country=Israel
[…]
organization=Department of Applied Mathematics and Scientific Computing, Indian Institute of Technology Roorkee,country=India
[…]
organization=Department of Mathematics, The Technion- Israel Institute of Technology,city=Haifa, country=Israel
1 Introduction
Let $\mathbb{R}^{m}$ denote the Euclidean space equipped with the usual inner product $\langle\cdot,\cdot\rangle$ and its associated norm $\|\cdot\|$. Let $f:\mathbb{R}^{m}\to\mathbb{R}$. Our goal is to solve the following minimization problem:
$\text{find }x^{*}\in\mathbb{R}^{m}\text{ such that }f(x^{*})=\min_{x\in\mathbb{R}^{m}}f(x).$ (1)
We denote the set of minimizers of $f$ by $\Omega_{\text{min}}(f,\mathbb{R}^{m})$.
A common approach to find a solution to (1) is the following iterative method: Let $x_{0}\in\mathbb{R}^{m}$ be an initial point. Given the current iterate $x_{n}$, the next iterate $x_{n+1}$ is computed by
$x_{n+1}=x_{n}+\alpha_{n}d_{n}\text{ for all }n\in\mathbb{N}_{0}=\mathbb{N}\cup\{0\},$ (2)
where $\alpha_{n}>0$ is the stepsize determined by exact or inexact line search and $d_{n}$ is a suitable search direction. Methods that follow the update rule (2) are referred to as line search methods. These methods require only the search direction $d_{n}\in\mathbb{R}^{m}$ and the stepsize $\alpha_{n}\in(0,\infty)$. An appropriate descent direction $d_{n}$ must be chosen so that it satisfies the descent condition:
$\langle\nabla f(x_{n}),d_{n}\rangle<0\text{ for all }n\in\mathbb{N}_{0}.$ (3)
A common choice for a descent direction is $d_{n}=-\nabla f(x_{n})$, which simplifies the general line search method (2) to the familiar gradient descent (GD) method:
$x_{n+1}=x_{n}-\alpha_{n}\nabla f(x_{n})\text{ for all }n\in\mathbb{N}_{0}.$
If there exists a constant $c>0$ such that
$\langle\nabla f(x_{n}),d_{n}\rangle\leq-c\|\nabla f(x_{n})\|^{2}\text{ for all }n\in\mathbb{N}_{0},$
then the vector $d_{n}$ is said to satisfy the sufficient descent condition.
Conjugate gradient (CG) methods are among the most effective algorithms for solving large-scale unconstrained optimization problems due to their low storage requirements. These methods require only first-order derivatives, makes them highly suitable for large-scale applications. The conjugate gradient (CG) method was originally proposed by Hestenes and Stiefel [Hestenes1952] for the solution of linear systems. Later, Fletcher and Reeves [Fletcher] extended this concept and formulated the first nonlinear CG method for unconstrained optimization. Due to their strong theoretical foundation and excellent numerical performance, many other versions of the CG method have been proposed (see [Polak1969, Dai2001, Fletcher1987, Dai1999] and the references therein).
In general, CG methods use the iterative scheme (2), where the search direction $d_{n}$ is defined by
$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\nabla f(x_{n})+\beta_{n-1}d_{n-1},&\text{ if }n\geq 1,\end{cases}$ (4)
where $\beta_{n}$ is the CG parameter. There are many formulas for $\beta_{n}$, each resulting in different computational behavior [HagerZhang2006].
One of the most well-known CG methods was proposed by Dai and Liao [Dai2001]. Their approach is based on an extended conjugacy condition, where the CG parameter is defined by
$\beta_{n}^{DL}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-t\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$ (5)
where $t\in(0,\infty)$, $s_{n}=x_{n+1}-x_{n}$, and $y_{n}=\nabla f(x_{n+1})-\nabla f(x_{n})$. Note that when $t=0$, $\beta_{n}^{DL}$ simplifies to the CG parameter proposed by Hestenes and Stiefel [Hestenes1952]:
$\beta_{n}^{HS}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
The CG parameter proposed by Hager and Zhang [HagerZhang2005] is defined by
$\beta_{n}^{HZ}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
The parameter $\beta_{n}^{HZ}$ can be viewed as an adaptive version of (5) for the specific choice $t=2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}$. Similarly, the CG parameter proposed by Dai and Kou [DaiKou2013] is defined by
$\beta_{n}^{DK}(\tau_{n})=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-\left(\tau_{n}+\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},y_{n}\rangle},$
where $\tau_{n}$ is a parameter that plays the role of a scaling factor in the scaled memoryless Broyden–Fletcher–Goldfarb–Shanno (BFGS) method, which can be viewed as another adaptive variant of
Motivated by Shanno’s matrix interpretation of CG methods [Shanno1978], Babaie-Kafaki and Ghanbari [KafakiGh2014] showed that the Dai–Liao search directions can be written in the form $d_{n+1}=-P_{n+1}\nabla f(x_{n+1})$ for all $n\in\mathbb{N}_{0}$, where $P_{n+1}=I-\frac{s_{n}y_{n}^{T}}{\langle s_{n},y_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}$ is a nonsingular matrix for $t\in(0,\infty)$ and $\langle s_{n},y_{n}\rangle\neq 0$; see also [KafakiEJOR2014]. By performing an eigenvalue analysis on the symmetrized version of $P_{n+1}$, defined by $A_{n+1}=\frac{P_{n+1}+P_{n+1}^{T}}{2}$, we can derive a two-parameter formula for $t$ in the Dai-Liao CG parameter (5), given by
$t_{n}^{p,q}=p\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}-q\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0},$ (6)
which ensures the descent condition (3) for $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. However, in cases where $\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}$ and $\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$ are nearly equal, their subtraction may not be numerically stable [Kafaki2016]. Moreover, the parameter $t$ in the CG parameter $\beta_{n}^{DL}$ must be nonnegative. To address these issues, they restricted $q$ in (6) to be nonpositive and proposed the following CG parameter:
$\beta_{n}^{DDL}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$
where $t_{n}^{p,q}$ is defined by (6) with $p\in\left(\frac{1}{4},\infty\right)$ and $q\in(-\infty,0]$. Note that for the parameter choice $(p,q)=(2,0)$ in (6), $\beta_{n}^{DDL}$ reduces to $\beta_{n}^{HZ}$. Similarly, if $(p,q)=(1,0)$, then $\beta_{n}^{DDL}$ is equivalent to $\beta_{n}^{DK}(\tau_{n})$ with the optimal choice of $\tau_{n}=\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$. For additional variants of the Dial–Liao conjugate gradient method, we refer to [Kafaki2025, Kafaki2025Opt].
Another significant category of iterative methods used for solving unconstrained optimization problems is the spectral gradient method. Initially introduced by Barzilai and Borwein [Barzilai1988], these methods were later advanced by Raydan [Raydon1997], who developed a spectral gradient approach specifically for large-scale unconstrained optimization challenges. The spectral conjugate direction is defined by
$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\theta_{n}\nabla f(x_{n})+\beta_{n-1}d_{n-1},&\text{ if }n\geq 1,\end{cases}$
where $\{\theta_{n}\}\subset(0,\infty)$ is the spectral parameter and $\{\beta_{n}\}\subset(0,\infty)$ is the conjugate parameter.
Building on the encouraging numerical results associated with spectral gradient methods [Barzilai1988, Raydon1997], Birgin and Martínez [Birgin2001] proposed the integration of spectral gradient techniques with conjugate gradient methods, culminating in the development of the first spectral conjugate gradient (CG) method. The spectral CG methods represent a powerful enhancement of traditional CG methods by incorporating the principles of spectral gradients within the conjugate gradient framework. Continuing in this direction, Yu and co-authors [Yu2008OMS] adapted Perry’s spectral CG formulation [Perry1978], obtaining a more effective variant of the method. Inspired by the good convergence of quasi-Newton techniques, Andrei [Andrei2007] proposed a CG method using a scaled BFGS preconditioner, and this idea was further developed in later works [Andrei2008, Andrei2010EJOR, Babaie2013MMA].
Li and Fukushima [Li2001] proposed a modified form of the secant condition, and Zhou and Zhang [Zhou2006] later built on their idea to obtain the following modified secant relation:
$\nabla^{2}f(x_{n+1})s_{n}=z_{n},$ (7)
where
$z_{n}=y_{n}+h_{n}\|\nabla f(x_{n})\|^{r}s_{n},\quad h_{n}=\nu+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|\nabla f(x_{n})\|^{-r}$ (8)
with $r,\nu\in(0,\infty)$. The modified secant condition (8) is essential to prove the global convergence of the MBFGS algorithm when dealing with non-convex objectives [Zhou2006]. Note that the following expression holds irrespective of the line search used.
$\langle z_{n},s_{n}\rangle$ $=\langle y_{n},s_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|^{2}+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|s_{n}\|^{2}$
$\geq\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|^{2}>0\text{ for all }n\in\mathbb{N}_{0}.$
Since (8) leads to favorable theoretical properties, Faramarzi and Amini [Faramarzi2019] modified the spectral CG method of Jian et al. [Jian2017OMS] and proposed the following choice for the conjugate parameter:
$\beta_{n}^{ZDK}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
Based on this conjugate parameter, they constructed the following spectral parameter:
$\theta_{n+1}=\begin{cases}1-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle\nabla f(x_{n+1}),z_{n}\rangle},&\text{ if }1-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle\nabla f(x_{n+1}),z_{n}\rangle}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise},\end{cases}$
where $\eta$ and $\tau$ are positive constants. To ensure global convergence of this method, Faramarzi and Amini [Faramarzi2019] used the line search that satisfied the strong Wolfe conditions
$f(x_{n}+\alpha_{n}d_{n})\leq f(x_{n})+\delta\alpha_{n}\langle\nabla f(x_{n}),d_{n}\rangle,$ (9)
$|\langle\nabla f(x_{n}+\alpha_{n}d_{n}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle,$ (10)
where $0<\delta<\sigma<1$, and the following assumptions:
  1. (A1)
    The level set $C_{0}=\{x\in\mathbb{R}^{n}:f(x)\leq f(x_{0})$ is bounded.
  2. (A2)
    $f$ is continuously differentiable in some neighbourhood $\Omega$ of $C_{0}$ with $L$-Lipschitz continuous gradient.
The primary motivation of the present study is to propose a modified spectral conjugate gradient method, guided by an eigenvalue-based analysis and a modified secant condition, and to prove global convergence for general objective functions under typical conditions.This new method always meets the sufficient-descent requirement, independent of the line search strategy applied. To demonstrate the effectiveness of our proposed approach, we present numerical results that validate its performance and show its superior convergence compared to the modified spectral conjugate gradient method of Faramarzi and Amini [Faramarzi2019], and the scaled conjugate gradient method of Mard and Fakhari [Mrad2024]. Furthermore, we explore its application to compressed sensing, emphasizing its practical relevance and significant advantages.
2 Preliminaries
In this section we present some fundamental results regarding Euclidean space that will be used to establish the convergence of our proposed method.
Lemma 1.
([Faramarzi2019]) Let $a,b,\sigma\in\mathbb{R}$. Then
$(c+\sigma d)^{2}\leq(1+\sigma^{2})(c^{2}+d^{2})\text{ for all }c,d,\sigma\geq 0.$
Lemma 2.
([Faramarzi2019]) Let $\mathbb{R}^{m}$ be a Euclidean space. Then
$\langle u,v\rangle\leq\frac{1}{2}(\|u\|^{2}+\|v\|^{2})\text{ for all }u,v\in\mathbb{R}^{m}.$
Lemma 3.
Let $\mathbb{R}^{m}$ be a Euclidean space, and let $\{u_{n}\}$ and $\{v_{n}\}$ be sequences in $\mathbb{R}^{m}$. Let $\sigma\in(0,\infty)$ and let $\{a_{n}\}$ and $\{b_{n}\}$ be two sequences in $(0,\infty)$ satisfying the following conditions:
  1. (C1)
    $a_{n}>a>0$ for all $n\in\mathbb{N}$,
  2. (C2)
    $\langle v_{n},u_{n}\rangle<0$ for all $n\in\mathbb{N}$,
  3. (C3)
    $|\langle v_{n+1},u_{n}\rangle|\leq-\sigma\langle v_{n},u_{n}\rangle$ for all $n\in\mathbb{N}$.
Suppose that
$u_{n+1}=-a_{n+1}v_{n+1}+b_{n}u_{n}\text{ for all }n\in\mathbb{N}.$ (11)
Then
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}\geq\frac{\|v_{n+1}\|^{4}}{\|u_{n+1}\|^{2}}\left[\frac{a^{2}}{1+\sigma^{2}}-a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\right].$
Proof.
Using (11) and condition (C2), we obtain
$b_{n}^{2}\|u_{n}\|^{2}$ $=\|u_{n+1}\|^{2}+2a_{n+1}\langle v_{n+1},u_{n+1}\rangle+a_{n+1}^{2}\|v_{n+1}\|^{2}$
$\leq\|u_{n+1}\|^{2}+a_{n+1}^{2}\|v_{n+1}\|^{2},$
which implies that
$\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\geq b_{n}^{2}-a_{n+1}^{2}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$ (12)
Again, using (11), we have
$\langle v_{n+1},u_{n+1}\rangle+b_{n}\langle v_{n+1},u_{n}\rangle=a_{n+1}\|v_{n+1}\|^{2}.$ (13)
It follows from condition (C3) that
$b_{n}\langle v_{n+1},u_{n}\rangle\leq b_{n}|\langle v_{n+1},u_{n}\rangle|\leq-\sigma b_{n}\langle v_{n},u_{n}\rangle\leq\sigma b_{n}|\langle v_{n},u_{n}\rangle|.$ (14)
Using (13) and (14), we get
$\langle v_{n+1},u_{n+1}\rangle+\sigma b_{n}|\langle u_{n},v_{n}\rangle|\geq a_{n+1}\|v_{n+1}\|^{2}.$
Squaring both sides and applying Lemma 1 with $c=-\langle v_{n+1},u_{n+1}\rangle>0$ (from (C2)) and $d=b_{n}|\langle v_{n},u_{n}\rangle|>0$ gives
$\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}\geq\frac{a_{n+1}^{2}}{1+\sigma^{2}}\|v_{n+1}\|^{4}.$
Since $a_{n+1}>a$, it follows that
$\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}$ $\geq\frac{a^{2}}{1+\sigma^{2}}\|v_{n+1}\|^{4},$
which implies that
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}$
$=\frac{1}{\|u_{n+1}\|^{2}}\left[\langle v_{n+1},u_{n+1}\rangle^{2}+\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\langle v_{n},u_{n}\rangle^{2}\right]$
$\geq\frac{1}{\|u_{n+1}\|^{2}}\left[\frac{a^{2}}{1+\sigma^{2}}\|v_{n+1}\|^{4}+\left(\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}-b_{n}^{2}\right)\langle v_{n},u_{n}\rangle^{2}\right].$ (15)
Using (12) and (15), we conclude that
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}\geq\frac{\|v_{n+1}\|^{4}}{\|u_{n+1}\|^{2}}\left[\frac{a^{2}}{1+\sigma^{2}}-a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\right].$
Lemma 4.
([Zoutendijk]) Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be such that assumptions (A1) and (A2) hold. For any iterative method of the form (2) with $d_{n}$ being a descent direction and $\alpha_{n}$ determined by the strong Wolfe conditions (9) and (10), we have
$\sum_{n=0}^{\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}<\infty.$
It follows from Lemma 4 that
$\lim_{n\to\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}=0.$ (16)
3 A modified descent Dai-Liao spectral conjugate gradient method and its convergence analysis
In this section we introduce a modified spectral conjugate gradient (CG) descent method, inspired by ideas from quasi-Newton methods.
Using the theoretical properties provided by (8) and the Dai-Liao conjugate gradient parameter (5), we propose the following modified Dai-Liao conjugate parameter:
$\beta_{n}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$ (17)
where $t$ is a nonnegative parameter.
Note that, by (4) and (17), the search direction can be expressed as
$d_{n+1}=-Q_{n+1}\nabla f(x_{n+1})\text{ for all }n\in\mathbb{N}_{0},$
where
$Q_{n+1}=I-\frac{s_{n}z_{n}^{T}}{\langle s_{n},z_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}.$
Following the methodology in [KafakiGh2014], we propose the following formula for the parameter $t$ in the modified Dai-Liao conjugate parameter (17):
$t_{n}^{p,q}=p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0}$ (18)
with $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. Using this parameter $t_{n}^{p,q}$, we define the following modified descent Dai-Liao (MDDL) conjugate parameter:
$\beta_{n}^{MDDL}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$ (19)
Now, we consider the following spectral conjugate gradient search direction:
$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\theta_{n}\nabla f(x_{n})+\beta_{n-1}^{MDDL}d_{n-1}&\text{ if }n\geq 1,\end{cases}$ (20)
where $\{\theta_{n}\}$ is a sequence in $(0,\infty)$.
Proposition 5.
Let $f:\mathbb{R}^{m}\to\mathbb{R}$ be a differentiable function and let $\{d_{n}\}$ be a sequence of spectral CG search directions generated by (20). Then the following statements hold:
  • (a)
    For all $n\in\mathbb{N}_{0},$ we have
    $\langle\nabla f(x_{n}),d_{n}\rangle\leq-\kappa_{n}\|\nabla f(x_{n})\|^{2},$ (21)
    where
    $\kappa_{n}=\left\{\begin{array}[]{ll}1,&\text{if }n=0,\\ \theta_{n}-\left(\frac{1}{4p}+|q|\right),&\text{if }n\geq 1.\end{array}\right$.
  • (b)
    If $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$, then the direction given by (20) is a descent direction.
Proof.
(a) Note that $\langle\nabla f(x_{0}),d_{0}\rangle=-\|\nabla f(x_{0})\|^{2}$. Therefore inequality (21) holds for $n=0$. We claim that the inequality (21) holds for $n>0$. Let $n\in\mathbb{N}_{0}$. Note that $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. For simplicity, let us denote $\nabla f(x_{n+1})=F_{n+1}$. Then, using (18), (19) and (20), we get
$\langle F_{n+1},d_{n+1}\rangle$
$=\langle F_{n+1},-\theta_{n+1}F_{n+1}+\beta_{n}^{MDDL}d_{n}\rangle$
$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\frac{t_{n}^{p,q}\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-\left(p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}.$
Since $s_{n}=\alpha_{n}d_{n}$, we have
$\langle F_{n+1},d_{n+1}\rangle$ $=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle}+q\frac{\langle d_{n},z_{n}\rangle}{\|d_{n}\|^{2}}\frac{\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}+q\frac{\langle d_{n},F_{n+1}\rangle^{2}}{\|d_{n}\|^{2}}.$ (22)
Using Lemma 2 with $u_{n}=\frac{1}{\sqrt{2p}}F_{n+1}$ and $v_{n}=\frac{\sqrt{2p}\langle F_{n+1},d_{n}\rangle z_{n}}{\langle d_{n},z_{n}\rangle},$ we obtain
$\frac{\langle F_{n+1},d_{n}\rangle\langle F_{n+1},z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\leq\frac{1}{4p}\|F_{n+1}\|^{2}+\frac{p\|z_{n}\|^{2}\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}.$ (23)
On the other hand, by the Cauchy-Schwarz inequality, we have
$q\frac{\langle d_{n},F_{n+1}\rangle^{2}}{\|d_{n}\|^{2}}\leq|q|\frac{\|d_{n}\|^{2}\|F_{n+1}\|^{2}}{\|d_{n}\|^{2}}=|q|\|F_{n+1}\|^{2}.$ (24)
Combining (22), (23) and (24), we conclude that
$\langle F_{n+1},d_{n+1}\rangle$ $\leq-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{1}{4p}\|F_{n+1}\|^{2}+|q|\|F_{n+1}\|^{2}$
$=-\left(\theta_{n+1}-\left(\frac{1}{4p}+|q|\right)\right)\|F_{n+1}\|^{2}.$
(b) Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$. Then, employing (21), we get $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Therefore, $\{d_{n}\}$ is a descent direction as claimed. ∎
Now, we construct the spectral parameter for our conjugate gradient method based on the conjugate parameter $\beta_{n}^{MDDL}$.
Quasi-Newton methods are a prominent class of iterative optimization methods that avoid computing the exact Hessian at every step. Their main goal is to retain the fast convergence associated with Newton’s method while keeping the computational burden low. Rather than evaluating the true Hessian, these methods update an approximate one at each iteration. Motivated by the structure of quasi-Newton directions, the search direction $d_{n}$ in (20) is constructed to emulate this behavior. Our objective, therefore, is to choose an appropriate spectral parameter $\theta_{n+1}$ such that
$B_{n+1}^{-1}\nabla f(x_{n+1})\approx-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n},$
where $B_{n+1}$ is an approximation of the Hessian matrix $\nabla^{2}f(x_{n+1})$. Taking the inner product with $B_{n+1}s_{n}$, we find that
$\langle s_{n},\nabla f(x_{n+1})\rangle\approx-\theta_{n+1}\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle,$
which implies that
$\theta_{n+1}\approx\frac{1}{\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle).$ (25)
From (7), $B_{n+1}$ must satisfy the modified secant condition $B_{n+1}s_{n}=z_{n}$. It follows from (25) that
${\theta_{n+1}}$ $\approx\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle z_{n},d_{n}\rangle)$
[…]
$\quad\left.\left.-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)\langle z_{n},d_{n}\rangle\right)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right$.
$\quad\left.-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)$
$=\frac{\langle s_{n},\nabla f(x_{n+1})\rangle+\langle\nabla f(x_{n+1}),z_{n}\rangle-t_{n}^{p,q}\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}$
$=1-\frac{{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}-\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}$ (26)
$=1-(t_{n}^{p,q}-1)\frac{\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}.$ (27)
Neglecting $\langle s_{n},\nabla f(x_{n+1})\rangle$ in (26), we arrive at
${\theta_{n+1}}=1-\frac{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}.$ (28)
We denote the $\theta_{n+1}$ obtained in equation (27) by $\theta_{n+1}^{R}$ and the $\theta_{n+1}$ obtained in equation (28) by $\theta_{n+1}^{N}$. To ensure that the sufficient-descent condition holds and to keep the spectral parameters bounded, we introduce the following choices for the spectral parameters:
$\theta_{n+1}=\begin{cases}\theta_{n+1}^{R},&\text{ if }\theta_{n+1}^{R}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise }\end{cases}$ (29)
or
$\theta_{n+1}=\begin{cases}\theta_{n+1}^{N},&\text{ if }\theta_{n+1}^{N}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise },\end{cases}$ (30)
where $\eta$ and $\tau$ are positive constants. Relations (29) and (30) lead to
$\frac{1}{4p}+|q|+\eta\leq\theta_{n+1}\leq\tau\textrm{ for all }n\in\mathbb{N}_{0}.$ (31)
Now, we propose the following modified descent Dai-Liao spectral conjugate gradient method (MDDLSCG):
Input: Let $x_{0}\in\mathbb{R}^{n}$ and let $\epsilon,\delta,\sigma,\eta,\tau,r,\nu\in(0,\infty)$ be such that
     $0<\delta<\sigma<1$.
Step 0: Set $d_{0}=-\nabla f(x_{0})$.
Step 1: If $\|\nabla f(x_{n})\|_{\infty}<\varepsilon$, stop, where $\|u\|_{\infty}=\max_{1\leq i\leq m}|u_{i}|$ for
    $u=(u_{1},\ldots,u_{m})\in\mathbb{R}^{m}$.
Step 2: Determine the step length $\alpha_{n}$ satisfying the strong Wolfe
    conditions (9) and (10).
Step 3: Set $x_{n+1}=x_{n}+\alpha_{n}d_{n}$.
Step 4: Compute $d_{n+1}$ by
$d_{n+1}=-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\text{ for all }n\in\mathbb{N}_{0},$
    where the CG parameter $\beta_{n}^{MDDL}$ is given by (19) and
      the spectral parameter $\theta_{n}$ is given by (29) or (30).
Step 5: Set $n=n+1$ and return to Step 1.
Algorithm 1 MDDLSCG Algorithm
Remark 6.
(i) Using $\theta_{n}\geq\frac{1}{4p}+|q|+\eta$ and (21), we conclude that
$\langle\nabla f(x_{n}),d_{n}\rangle\leq-\eta\|\nabla f(x_{n})\|^{2}\text{ for all }n\in\mathbb{N}_{0}.$ (32)
(ii) If $\|\nabla f(x_{n})\|\neq 0$, then using (10) and (32), we see that
$\langle d_{n},y_{n}\rangle\geq-(1-\sigma)\langle\nabla f(x_{n}),d_{n}\rangle\geq(1-\sigma)\eta\|\nabla f(x_{n})\|^{2}>0,$ (33)
which, by (8), implies that $h_{n}=\nu$ and hence
$z_{n}=y_{n}+\nu\|\nabla f(x_{n})\|^{r}s_{n}.$ (34)
Now, we analyze the global convergence of the MDDLSCG Algorithm.
Remark 7.
  1. (i)
    Employing (9) and (32), we can see that $\{f(x_{n})\}$ is decreasing. Thus, the sequence $\{x_{n}\}$ generated by Algorithm 1 is contained in $C_{0}$.
  2. (ii)
    From Assumption (A1), there exists $M>0$ such that $\|x\|\leq M$ for all $x\in C_{0}$. Then using Assumption (A2), we have
    $\|\nabla f(x)\|$ $\leq\|\nabla f(x)-\nabla f(x_{0})\|+\|\nabla f(x_{0})\|$
    $\leq L\|x-x_{0}\|+\|\nabla f(x_{0})\|$
    $\leq 2ML+\|\nabla f(x_{0})\|=\gamma\text{ for all }x\in C_{0}.$ (35)
In addition, suppose that $\nabla f(x_{n})\neq 0$ for all $n\in\mathbb{N}_{0}$; otherwise, a stationary point will be found.
Lemma 8.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$. If $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$, then $\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}}<\infty.$
Proof.
Since $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$, there exists $\epsilon>0$ such that $\|\nabla f(x_{n})\|\geq\epsilon$ for all $n\in\mathbb{N}_{0}$. When combined with (16), this implies that
$\lim_{n\to\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}}=0.$ (36)
Let $u_{n}=d_{n}$, $v_{n}=\nabla f(x_{n})$, $a_{n}=\theta_{n}$ and $b_{n}=\beta_{n}$ in Lemma 3. Since $\theta_{n}>\frac{1}{4p}+|q|>0$, it follows from Remark 6 that $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Also from strong Wolfe condition (10), we have $|\langle\nabla f(x_{n+1}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle$ for all $n\in\mathbb{N}_{0}$. Then, clearly all the conditions (C1), (C2) and (C3) of Lemma 3 are satisfied. Thus, we have
$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}\left[B-\theta_{n+1}^{2}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}}\right],$
where $B=\left(\frac{1}{4p}+|q|\right)^{2}\frac{1}{1+\sigma^{2}}$. It follows from (36) that for all sufficiently large $n$, we have
$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq B\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}.$ (37)
Using (37) and Lemma 4, we get that $\sum_{n=1}^{\infty}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}<\infty.$
Theorem 9.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a real function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Then $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0.$
Proof.
Suppose, for the sake of contradiction, that the assertion of the theorem does not hold. Then there exists a positive constant $\xi$ such that
$\|\nabla f(x_{n})\|>\xi\text{ for all }n\in\mathbb{N}_{0}.$ (38)
Employing (33), (34) and (38), we obtain
$\langle d_{n},z_{n}\rangle$ $=\langle d_{n},y_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}$
$\geq\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}\geq\nu\xi^{r}\alpha_{n}\|d_{n}\|^{2}.$ (39)
Using Assumption (A2), (34) and ((ii)), we have
$\|z_{n}\|\leq\|y_{n}\|+\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|\leq L\|s_{n}\|+\nu\gamma^{r}\|s_{n}\|=(L+\nu\gamma^{r})\|s_{n}\|.$ (40)
Thus, relations (19), ((ii)), (39), (40), along with the Cauchy-Schwarz inequality, yield
$\|\beta_{n}^{MDDL}\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\left(p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|+p\left\|\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$\quad+|q|\left\|\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\frac{\|\nabla f(x_{n+1})\|\|z_{n}\|}{\langle d_{n},z_{n}\rangle}+p\frac{\|z_{n}\|^{2}\|\nabla f(x_{n+1})\|\|d_{n}\|}{\langle d_{n},z_{n}\rangle^{2}}+|q|\frac{\|\nabla f(x_{n+1})\|}{\|d_{n}\|}$ (41)
$\leq\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}\|d_{n}\|}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}\|d_{n}\|}+|q|\frac{\gamma}{\|d_{n}\|}.$
Note that $\theta_{n+1}\leq\tau$ by (31). Hence, using (20) and (41), we arrive at
$\|d_{n+1}\|$ $=\|-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\|$
$\leq\|\theta_{n+1}\|\|\nabla f(x_{n+1})\|+\|\beta_{n}^{MDDL}\|\|d_{n}\|$
$\leq\tau\gamma+\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}}+|q|\gamma.$ (42)
Using (42), we observe that
$\frac{1}{\|d_{n+1}\|^{2}}\geq\frac{1}{E^{2}}\textrm{ for all }n\in\mathbb{N}_{0},$
where $E=\tau\gamma+\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}}+|q|\gamma$. Thus,
$\sum_{n=0}^{\infty}\frac{1}{\|d_{n}\|^{2}}=\infty.$ (43)
Employing (38) and Lemma 8, we find that
$\sum_{n=0}^{\infty}\frac{1}{\|d_{n}\|^{2}}=\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}\|\nabla f(x_{n})\|^{4}}\leq\frac{1}{\xi^{4}}\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}}<\infty,$
which contradicts (43). Therefore, $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0$, as asserted. ∎
4 Numerical Experiments
In this section we present the numerical performance of our proposed method (MDDLSCG), comparing it with Algorithm 3.1 in [Faramarzi2019] (denoted by MSCG) and the scaled conjugate gradient method (denoted by ScCG) in [Mrad2024]. All numerical experiments were conducted in MATLAB R2024b on an HP workstation with Intel Core i7-13700 processor, 32 GB RAM and Windows 11.
For the numerical experiments, we consider two test problems: (i) finding the minima of the Beale function, and (ii) finding the minima of the Almost Perturbed Quadratic (APQ) function.
Beale Test Function: The Beale function is multimodal with sharp peaks at the corners of the input domain, which is defined from the general Beale function as follows:
$f(x,y)=\left(1.5-x+xy\right)^{2}+\left(2.25-x+xy^{2}\right)^{2}+\left(2.625-x+xy^{3}\right)^{2}\text{ for all }(x,y)\in\mathbb{R}^{2}.$
The global minimum is located at $(x,y)=(3,0.5)$, where $f(x,y)=0$.
(a) Surface and Contour plot of the Beale functions
For the numerical tests, we choose $\delta=0.01$, $\sigma\in\{0.1,0.2,0.4,0.6,0.8,0.9\}$ for all the algorithms, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$ for algorithms MSCG and MDDLSCG and $p=0.4$ and $q=0.2$ for the algorithm MDDLSCG. We take the initial point $x_{0}=(1,0.8)$. The iterations are terminated once they satisfy the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-15}$.
The numerical results of all the algorithms MDDLSCG, MSCG and ScCG for the Beale function are summarized in Table 1, while their performance is illustrated in Figure 8(a) and 8(b). In Figures 8(a) (a), (c), (e), the iteration paths are visualized on the contour plot of the Beale function, showing how each method converges to the minimizer for $\sigma=0.1,0.2,0.6$, respectively. Figures 8(a) (b), (d), (f) depict the convergence rate by plotting the norm of the gradient, $\|\nabla f(x_{n})\|_{\infty}$ versus the number of iterations on a logarithmic scale for $\sigma=0.1,0.2,0.6$, respectively. The comparison graph between different values of sigma and the number of iterations is depicted in Figure 8(b) (a) and the comparison graph between different values of sigma and execution time is in Figure 8(b) (b).
In view of the data presented, it is evident that our proposed method, MDDLSCG, outperforms both MSCG and ScCG in all tested scenarios. As shown in Table 1, Figure 8(a) and 8(b) (b), MDDLSCG reaches the stopping criteria in fewer iterations and requires less execution time. Additionally, using Figure 8(b) (a), we observe that these methods are significantly sensitive to variations in the parameter $\sigma$, which affects their performance.
$\sigma$ MDDLSCG MSCG ScCG itr Tcpu $E_{n}$ itr Tcpu $E_{n}$ itr Tcpu $E_{n}$ 0.1 37 1.86e-02 9.49e-15 63 3.14e-02 2.63e-16 200 5.40e-02 5.27e-16 0.2 36 1.75e-02 3.16e-15 63 2.76e-02 9.07e-15 136 3.99e-02 4.51e-15 0.4 53 2.44e-02 9.40e-15 73 2.65e-02 8.75e-15 126 4.73e-02 2.87e-15 0.6 40 2.36e-02 7.91e-16 72 2.78e-02 5.74e-15 95 4.94e-02 9.32e-15 0.8 67 3.24e-02 5.32e-15 111 2.86e-02 4.48e-15 201 4.92e-02 6.91e-15 0.9 64 3.28e-02 8.16e-15 101 3.48e-02 8.74e-15 184 4.37e-02 1.08e-15
Table 1: Comparison table for MDDLSCG, MSCG and ScCG for Beale function
(a) Comparison of the iterations and execution time for different $\sigma$ for MDDLSCG, MSCG, ScCG for the Beale function
(b) Comparison of the iterations and execution time for different $\sigma$ for MDDLSCG, MSCG, ScCG for the Beale function
Almost Perturbed Quadratic function: Consider the Almost Perturbed Quadratic (APQ) function $f:\mathbb{R}^{m}\to\mathbb{R}$ defined by
$f(x)=\frac{1}{2}x^{T}Ax+b^{T}x\text{ for all }x\in\mathbb{R}^{m},$
where $A\in\mathbb{R}^{m\times m}$ is a symmetric positive definite matrix ans $b\in\mathbb{R}^{m}$ is a vector of perturbations. This function has a global minimum at $x=-A^{-1}b$. Note that the gradient of $f$ is given by
$\nabla f(x)=Ax+b\text{ for all }x\in\mathbb{R}^{m}.$
For the numerical experiments, we have considered three cases, each characterized by dimensions of $1000$, 100100 and $25$. In the cases with dimensions $1000$ and $100$, the matrix $A$ and vector $b$ are randomly generated using MATLAB’s built-in function rand and the initial point is randomly generated using MATLAB’s rand function.
In the case of dimension $25$, we take the initial point
$x_{0}$ $=(x_{0,1},\ldots,x_{0,i},\ldots,x_{0,25}),\text{ where }x_{0,i}=1\text{ for all }i=1,2,\ldots,25\text{ and}$
$a$ $=(2,4,3,6,4,12,8,10,5,8,6,11,4,8,3,6,2,12,6,8,4,10,5,2,6),$
$b$ $=(-4,-52,-18,-60,-24,-72,-72,-50,-55,-80,-78,-22,-12,$
$\quad-80,-27,-42,-24,-120,-30,-120,-24,-100,-60,-20,-66),$
$A$ $=\text{diag}(a).$
In all cases, the parameters are set as follows: $\delta=0.01$ and $\sigma=0.1$ for all algorithms, while for MSCG and MDDLSCG, we use $\eta=0.001$, $\tau=10$, $r=1$, and $\nu=0.001$. Additionally, for the MDDLSCG algorithm, we set $p=0.4$ and $q=0.2$.
The iterations are terminated once they satisfy the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-6}$ is satisfied, where $x_{n}=(x_{n,1},\ldots,x_{n,25})$.
The performance of the algorithms MDDLSCG, MSCG, and ScCG for the APQ function is summarized in Table 2 for all three cases. To provide a clearer visualization of the experiment, we have plotted the graph of the number of iterations versus $E_{n}$ in Figure 11(a). From Table 2 and Figure 11(a), it is evident that our proposed method, MDDLSCG, requires fewer iterations and less computational time to satisfy the desired stopping criterion across all cases.
Figure 11(b) illustrates how the algorithms MDDLSCG, MSCG, and ScCG approximate the desired solution $x^{*}=-A^{-1}b$ at the 5th and 10th iterations in the case where the problem dimension is $25$.
Furthermore, Figure 11(c) depicts the convergence behavior of the algorithms as the number of iterations increases up to the 50th iteration for the same dimension. It is observed that MDDLSCG demonstrates superior convergence characteristics compared to MSCG and ScCG.
Dimension Method itr Tcpu $Er_{n}$
25 MDDLSCG 52 3.088055e-02 5.691882e-07
MSCG 59 3.947327e-02 6.931169e-07
ScCG 65 4.730704e-02 4.514777e-07
100 MDDLSCG 102 5.552326e-02 8.162825e-07
MSCG 121 6.512598e-02 9.411660e-07
ScCG 134 6.823947e-02 8.609522e-07
1000 MDDLSCG 132 3.481332e-02 9.291368e-07
MSCG 146 3.522395e-02 9.484382e-07
ScCG 171 4.053286e-02 8.747078e-07
Table 2: Comparison table of MDDLSCG, MSCG and ScCG for the APQ function
(a) Convergence behavior of $\|\nabla f(x_{n})\|_{\infty}$ for MDDLSCG, MSCG, ScCG for the APQ function
(b) Tracking behaviour of MDDLSCG, MSCG, ScCG for the APQ function at various iterations
(c) Graphical progression of MDDLSCG, MSCG, ScCG upto 50th iterations for the APQ function
5 Application to compressed sensing
Compressed sensing is a widely recognized signal processing approach that enables the efficient acquisition and reconstruction of signals with a sparse structure, allowing these signals to be stored in a compact and memoryless form [Bruckstein2009]. Its practical significance has been demonstrated in numerous fields, including machine learning, compressive imaging, radar systems, wireless sensor networks, medical and astronomical signal processing, and video compression, as discussed in [Aminifard2022].
The compressed sensing problem involves finding a sparse solution to a highly underdetermined system, represented as $Ax=b$, where $A\in\mathbb{R}^{m\times n}$ (with $m\ll n$) and $b\in\mathbb{R}^{m}$ [Bruckstein2009]. A common approach is to solve the following optimization problem:
$f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\phi(x),$
where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a penalty function, and $\mu>0$ is a parameter that regulates the trade-off between sparsity and reconstruction fidelity. A popular choice for $\phi$ is $\phi(x)=\|x\|_{1}$, leading to the well-known Basis Pursuit Denoising (BPD) problem, which has been extensively studied in the literature; see [Esmaeili2019] and references therein. However, solving BPD is difficult because the $\ell_{1}$-penalty term is non-smooth. To overcome this difficulty, Zhu et al. [Zhu2013] introduced a relaxed form of BPD based on Nesterov’s smoothing technique [Nesterov2005Siam], leading to the following formulation:
$f_{\lambda}(x)=\sum_{i=1}^{n}\psi_{\lambda}(|x_{i}|),$
where $x_{i}$ represents the $i$-th component of the vector $x$, and $\psi_{\lambda}$ is defined by
$\psi_{\lambda}(|x_{i}|)=\max_{\gamma\in[0,1]}\left\{\gamma|x_{i}|-\frac{1}{2}\lambda\gamma^{2}\right\},$
which simplifies to:
$\psi_{\lambda}(|x_{i}|)=\begin{cases}\frac{|x_{i}|^{2}}{2\lambda},&\text{if }|x_{i}|<\lambda,\\ |x_{i}|-\frac{\lambda}{2},&\text{if }|x_{i}|\geq\lambda,\end{cases}$
where $\lambda>0$. Function $f_{\lambda}$ is known as the Huber function [Rangarajan], and its gradient $\nabla f_{\lambda}$ is $\frac{1}{\lambda}$ -Lipschitz continuous [Becker2011]. The explicit form of the Huber function is given by
[…]
where $\text{sgn}(\cdot)$ denotes the signum function.
Therefore, instead of solving the original problem for $f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\|x\|_{1}$, we solve the smoothed version:
$\min_{x\in\mathbb{R}^{n}}f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu f_{\lambda}(x).$
Clearly, the problem above is an unconstrained smooth convex optimization problem, and its gradient is given by:
$\nabla f(x)=\lambda\nabla f_{\lambda}(x)+A^{T}(Ax-b).$
For the numerical experiments, we aim to reconstruct a sparse signal of length $n$ from $m$-length observations ($m\ll n$). The parameters for the MDDLSCG, MSCG and ScCG algorithms are set as follows:
$\mu=\max\{2^{-7},0.001\|A^{T}b\|_{\infty}\}$, $\lambda=\min\{0.001,0.048\|A^{T}b\|_{\infty}\}$, $\sigma=0.1$, $\delta=0.01$, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$.
For MDDLSCG, we choose $p=0.4$ and $q=0.2$. The Gaussian matrix $A$ is generated with the MATLAB commands randn(m, n) and the noise vector $w$ is generated with $0.1\times\texttt{randn(m, 1)}$. The measurement vector $b$ is computed by $b=Ax+w.$ The regularization parameter is chosen as $\mu=0.001\|A^{T}b\|_{\infty}$. Each methods begin with the initial point $x_{0}=A^{T}b$. Let $x$ represent the true signal and $x^{*}$ denote the recovered one. The termination criterion for the experiment is the mean-squared error(MSE):
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-x_{i}^{*})^{2}\leq 10^{-5}.$
To assess the reconstruction quality, we also compute the relative error (RelErr) of the restored signal, defined by
$\text{RelErr}=\frac{\|x^{*}-x\|}{\|x\|}.$
Let $k$ represent the number of nonzero components in the original signal. To ensure a fair comparison, we test two cases: $(m,n,k)=(128,512,16)$, $(256,1024,32)$.
The numerical results are summarized in Table 3, where we list, for each case, the average iteration count (Itr), the average CPU time (Tcpu), the mean squared error (MSE), and the relative error (RelErr). The graph depicting the relationship between relative error and the number of iterations is shown in Figures 11(d)(a) and 13(a)(a), while the graph illustrating MSE versus the number of iterations is presented in Figures 11(d)(b) and 13(a)(b). In addition, the recovery of the signal is visualized in Figures 11 and 13.
From Table 3, Figures 11(d) and 13(a), it is clear that the MDDLSCG algorithm surpasses the MSCG and ScCG algorithms in performance across all cases.
(d) Convergence behavior and relative error plots, demonstrating the performance of the methods MDDLSCG, MSCG and ScCG for $(m,n,k)=(128,512,16)$
Figure 11: Signal recovery for data $(m,n,k)=(128,512,16)$
(a) Convergence behavior and relative error plots, demonstrating the performance of the methods MDDLSCG, MSCG and ScCG for $(m,n,k)=(256,1024,32)$
Figure 13: Signal recovery for data $(m,n,k)=(256,1024,32)$
Dim(m,n,k)(m,n,k) Method Itr Tcpu MSE RelErr (128,512,16) MDDLSCG 272 1.186382e-01 2.033486e-06 7.487872e-03 MSCG 493 1.606987e-01 2.033495e-06 7.488087e-03 ScCG 3653 1.135658e+00 2.033588e-06 7.712389e-03 (256,1024,32) MDDLSCG 291 2.163503e-01 2.128598e-06 9.721181e-03 MSCG 1405 8.339790e-01 2.128516e-06 9.720993e-03 ScCG 3002 1.746555e+00 2.229608e-04 9.949161e-02
Table 3: Comparison table for methods MDDLSCG, MSCG and ScCG
6 Conclusions
We investigated a modified descent Dai-Liao spectral conjugate gradient (MDDLSCG) method that incorporates a new secant condition and quasi-Newton direction. The updated method introduces spectral parameters ensuring a sufficient descent property independent of line search techniques. Theoretical analysis establishes the global convergence of our method for general nonlinear functions under standard assumptions. Numerical experiments and an application to signal processing demonstrate the improved performance and practical effectiveness of the proposed algorithm. Additionally, we present numerical results that validate its performance and show its superior convergence compared to the MSCG method [Faramarzi2019] and the ScCG method [Mrad2024], highlighting its potential in optimization and related applications.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding
The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.